The area of a trapezium is 364-5 sq.m.Its parallel sides are in the ratio 5:4.If perpendicular distance between parallel lines (sides) is 27m, then the bigger parallel side will be:
Answer by mananth(16946)

Area of trapezium = 364.5
let smaller side be x
and bigger side be y
x/y = 5/4
let the common multiple be k
x=5k
y=4k
Area = (1/2)(5k+4k)*27
364.5=(1/2)(5k+4k)*27
multiply by 2
729=(5k+4k)*27
/27
729/27 =9k
729/(27*9) =k
k=3
larger side = 5*3=15m